OpenStudy (anonymous):

Which of the following is the simplified form of ? (x^2-9)/(x-3)

OpenStudy (anonymous):

x+3

OpenStudy (asnaseer):

remember that:\[a^2-b^2=(a+b)(a-b)\]so:\[x^2-9=x^2-3^2=(x+3)(x-3)\]this can be used to cancel the (x-3) terms from numerator and denominator to get the answer @jimmyrep gave.

OpenStudy (anonymous):

A,b,c or d x – 3, with the restriction x ≠ 3 x + 3, with the restriction x ≠ 3 x – 3, with the restriction x ≠ –3 x + 3, with the restriction x ≠ –3

OpenStudy (asnaseer):

remember we got the answer x+3 by dividing the numerator and denominator by (x-3). we cannot do this if x=3 as then we would be dividing by zero which is not a valid operation.

OpenStudy (asnaseer):

so what do you think the answer is?

OpenStudy (anonymous):

i said b

OpenStudy (asnaseer):

perfect! - well done.

OpenStudy (anonymous):

thanks